Thinking About International Capital Flows

David Backus, Thomas Cooley, and Espen Henriksen Federal Reserve Bank of Kansas City May 21, 2012

This version: October 9, 2012

Backus, Cooley, and Henriksen (NYU)  International Capital Flows  1 / 36

Plan of attack

Thoughts about capital flows Evidence Demography Parts of a model Speculating about China

Thoughts

Thoughts

Backus, Cooley, and Henriksen (NYU)  International Capital Flows  2 / 36

Thoughts

Thoughts about capital flows

Michael Bordo, “Globalization in historical perspective,” 2002

The fifty years before World War I saw massive flows of capital from Western Europe to (mainly) the Americas and Australasia. At its peak, the outflow from Britain reached nine percent of GNP and was almost as high in France, Germany, and the Netherlands. Private capital moved essentially without restriction ... into bonds that financed railroads and other infrastructure — and into government debt.

Link: http://econweb.rutgers.edu/bordo/nabe.pdf

Thoughts about capital flows

Lars Jonung, “Sweden under the gold standard,” 1984

In the second half of the nineteenth century, Sweden began a process of industrialization financed by considerable capital imports which transformed a basically agrarian country into an industrialized society.

From Bordo & Schwartz, Retrospective on the Gold Standard

Thoughts about capital flows

Michael Bordo, “Globalization in historical perspective,” 2002

A striking feature is the size and persistence of current account deficits in this period, esp in Australia, Canada, Argentina, and the Nordic countries, as well as the surpluses of the UK and France.

This globalization of finance had its dark side — in periodic crises when capital flows abruptly reversed themselves.

Link: http://econweb.rutgers.edu/bordo/nabe.pdf

Thoughts about capital flows

John Maynard Keynes, 1941-42

I disagree most strongly that the control of capital movements may be unnecessary. ... I see no reason to feel confidence that the more stable conditions [of the post-war era] will remove the more dangerous movements [of capital].

Central control of capital movements, both inward and outward, should be a permanent feature of the post-war system.

Quoted by Crotty, JEL, 1983

Thoughts about capital flows

Daniel Gross, New York Times, May 8, 2005

[US] imbalances are eerily reminiscent of recent economic crises. Could we see a perfect storm [for the US economy]? If so, what would it look like?

⋆ Nouriel Roubini estimates that long-term interest rates in the US could rise sharply and the dollar fall.

⋆ Jeffrey Frankel adds: “some of us have been warning of this hard-landing scenario for more than 20 years.”

Thoughts about capital flows

Øystein Olsen, Norges Bank, March 2011

Global trade imbalances have been reduced somewhat over the past two years, but there is a considerable risk that they will persist. They must be corrected.

Group of 20, Communique, April 15-16, 2011

We agreed on a set of indicative guidelines ... to address persistently large imbalances. We now launch ... an in-depth assessment of the nature of these imbalances and the root causes of impediments to adjustment. ... We will ascertain for our next meeting the corrective and preventive measures.

Thoughts about capital flows

Dick Caves and Ron Jones, World Trade and Payments, 1973

By the accounting rule, a preponderance of credits (or receipts) makes the balance positive or “favorable.” A balance with a preponderance of debits is called “unfavorable.” Notice the gravitational pull of semantics! The receipts side owns all the good words, and has done so since the eighteenth-century mercantilists made a virtue of “storing up treasure.”

Thoughts

Thoughts: summary

Long history of debate about capital flows

▶ Facilitator of economic growth? ▶ Or source of instability?

Suggestion: don’t let the words decide

▶ Global find and replace:

“global imbalances” → “international capital flows”

Evidence

Evidence

Backus, Cooley, and Henriksen (NYU)  International Capital Flows  11 / 36

Evidence

Evidence: current accounts

Largest deficits Amount (b$) Largest surpluses Amount (b$)
United States  –470 China  306
Italy  –72 Japan  195
Spain  –63 Germany  176
United Kingdom  –56 Mideast oil  150
France  –53 Switzerland  75
India  –49 Russia  71
Turkey  –49 Netherlands  56
Canada  –49 Norway  53
Brazil  –47 Taiwan  41
Greece  –32 Sweden  30
Australia  –32 Korea  28
Portugal  –23 Malaysia  28
Total deficits  –1183 Total surpluses  1466

Source: IMF, WEO, April 2011 version, data for 2010

Backus, Cooley, and Henriksen (NYU)  International Capital Flows  12 / 36

Evidence

Evidence: current accounts

Evidence

Evidence: net foreign assets

Evidence

Evidence: capital flows and oil

Evidence

Evidence: current account spectra

Evidence

Evidence: summary for capital flows and stocks

Capital flow variation is low-frequency [so much for HP filtering] Ditto net foreign assets Needed: an explanation for low-frequency movements

Demography

Demography

Backus, Cooley, and Henriksen (NYU)  International Capital Flows  18 / 36

Demography

Evidence: demography

Demography

Evidence: summary

Capital flows are persistent

▶ Ditto net foreign assets

Demography inherently persistent

▶ With sharp differences across countries

Worth exploring a connection? [Apparently yes: Attanasio-Kitao-Violante, Bloom-Canning, Boersch-Supan-Winter, Brooks, Domeij-Floden, Feroli, Ferrero, Henriksen, Krueger-Ludwig, and others had the same idea]

Model

Model

Backus, Cooley, and Henriksen (NYU)  International Capital Flows  21 / 36

Model

Model: mechanics

One-good deterministic world Countries: G7 (for now) Each has OG structure Power utility, fixed labor supply [+ age-specific adjustments] Identical production functions, country-specific productivity Fixed retirement age: 65 No bequests: perfect insurance for mortality risk No frictions either within or across countries

Model

Model: how it works

NFA is difference between net worth and capital stock

▶ Think of each as a function of age ▶ Aggregates are sums across age groups

Net worth follows from life-cycle saving

▶ Net worth varies by age ▶ Aggregate net worth depends on age distribution

Capital allocated to equate marginal products

▶ Connected to age distribution via effective labor input ▶ Equates K/Y across countries

Model

Model: questions

Individual saving rates differ across countries

▶ Not clear why

▶ Should we add “wedges” to mimic?

Capital-output ratios differ across countries

▶ Ditto

Market value ≠ capital stock

▶ Ditto

Model

Model: parameter values

Time interval  one year
Capital share (α)  1/3
Risk aversion (σ)  4
Discount factor (β)  1/1.025
Depreciation rate (δ)  2.5%
Fertility and mortality  UN data
Cohort sizes  UN data

Backus, Cooley, and Henriksen (NYU)  International Capital Flows  25 / 36

Model

Model: net foreign assets

China

China

Backus, Cooley, and Henriksen (NYU)  International Capital Flows  27 / 36

China

What about China?

Investment

▶ Lots of investment ▶ But capital-output ratio seems fine ▶ What’s missing?

Saving

▶ We can explain a high saving rate ▶ But not as high as we see ▶ What’s missing?

China

Why are China’s saving and investment so high?

China

Does China have too much capital?

No! Capital-output ratio not unusual Why not?

▶ PWT implies part of investment is high prices ▶ Growth central to steady state capital-output ratio

K/Y = \frac{I /Y}{g + δ}

[When you grow this fast, you need a lot of investment]

China

Does China have too little capital?

Stories — and a look around — suggest lots of capital projects Esp public infrastructure Also shortage of funds for private firms Can both be true?

China

Why is saving so high?

Demography gets us only part way Precautionary saving? High “business saving”?

Is China like Norway?

Norwegian Petroleum Fund (“Pension Fund Global”)

▶ Oil royalties not income to households ▶ Invested abroad by government

Story about China

▶ Business saving primarily SOEs ▶ Not income to households ▶ Invested abroad by government ▶ Like Norway?

China

Is convertibility an issue?

Category  Assets (tr$) Liabilities (tr$)
Direct investment  0.230  0.997
Portfolio investment  0.243  0.190
Other  0.535  0.451
Official reserves  2.453  —
Total  3.460  1.638
Net foreign assets  1.822

Source: IMF, IFS, May 2011 version, data for 2009

Final thoughts

Last thoughts

Capital flows

▶ A fact of life ▶ Persistent ▶ A role for demography? ▶ What else would you suggest?

Open questions

▶ Why are capital-output ratios so different across countries? ▶ Why are saving rates so different across countries? ▶ Why is China’s so high?